40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 8 16:27:13 2009, Jenne, Update, COC, ITMX wiped 
    Reply  Wed Dec 9 18:01:06 2009, Koji, Update, COC, Wiping finished 
Message ID: 2367     Entry time: Tue Dec 8 16:27:13 2009     Reply to this: 2373
Author: Jenne 
Type: Update 
Category: COC 
Subject: ITMX wiped 

Jenne, Kiwamu, Koji, Alberto, Steve, Bob

ITMX was wiped without having to move it. 
After 'practice' this morning on ETMY, Kiwamu and I successfully wiped ITMX by leaning into the chamber to get at the front face. 

Most notable (other than the not moving it) was that inspection with the fiber light before touching showed many very small particles on the coated part of the optic (this is versus ETMY, where we saw very few, but larger particles).  The after-wiping fiber light inspection showed many, many fewer particles on the optical surface.  I have high hopes for lower optical loss here!

ELOG V3.1.3-